OpenStudy (anonymous):

Simply the expression: 5^2 + 5^3 - 5^4

OpenStudy (anonymous):

5^2=25 +5^3=125-5^4=625=-475

OpenStudy (anonymous):

\[5^2 + 5^3 - 5^4 =5^2+5^2.5-5^2.5^2=5^2(...)\]

OpenStudy (anonymous):

5^2=25 +5^3=125-5^4=625=-475